OpenStudy (anonymous):

Write as a single logarithm 3loga x + 4loga y- 1/3loga z

OpenStudy (raden):

use the properties of logarithm bellow : * m loga c = loga (c^m) ** loga b + loga c = loga (b*c) *** loga b - loga c = loga (b/c)

OpenStudy (anonymous):

Sorry I'm do far beyond confused with this problem

OpenStudy (anonymous):

\[3\log_a x + 4\log_a y- \frac{1}{3}\log_a z=\log_a x^3 + \log_a y^4- \log_a z^{\frac{1}{3}}\]Power rule first.

OpenStudy (anonymous):

Then the product and quotient rules\[=\log_a \frac{x^3y^4}{z^{\frac{1}{3}}}\]